OpenStudy (anonymous):

What is the exact value of sqrt 38 / 2 sqrt 6

OpenStudy (anonymous):

Is the following the problem expression?\[\frac{\sqrt{38}}{2\sqrt{6}} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{\sqrt{38}}{2\sqrt{6}}\text{ = }\frac{\sqrt{19}\sqrt{2}}{2\sqrt{2}\sqrt{3}}\text{ = }\frac{\sqrt{\frac{19}{3}}}{2}\]